# DP-RDM: Adapting Diffusion Models to Private Domains Without Fine-Tuning

### v1

@@ 2.2 Differential Privacy
[…]
@@ 4.2 Quantitative Results
We first demonstrate the performance of DP-RDM quantitatively in terms of the Fréchet inception distance (FID) score. Figure 6 shows the FID score for DP-RDM on MS-COCO and Shutterstock when answering $T$ queries privately for $T$ ranging from 1 to 10k. The evaluation dataset and the private retrieval dataset are the same, i.e. for Figure 6(a) we treat the MS-COCO validation set as the evaluation dataset and MS-COCO training set as the private retrieval dataset. We highlight two notable observations.
@@ 4.3 Qualitative Results
Ref
Validation Prompt
1
image of blue sky with white cloud for background usage.
2
wood texture paint yellow
3
School of Jackfish in Sipadan Malaysia
4
A bunch of garlic isolated on white background
5
Halloween pumpkins on wooden table on dark color background
6
cauliflower
7
Closeup of golden art sculpture pagoda and golden tiered umbrella in temple against with blue sky in north of Thailand. Faith and religion concept.
8
White stork hunting in a meadow
9
abstract light background
10
White grunge wood close-up background. Painted old wooden wall.
11
Multicolored abstract background holidays lights in motion blur image. illustration digital.
12
Torn blue jean fabric texture.
@@
Ref
Validation Prompt
1
A dog watching a television that is displaying a picture of a dog.
2
The Asian girls are sharing conversation under a parasol.
3
An aerial view of several jumbled cars on a narrow road.
4
a short passenger train on a track by a grassy hill
5
A kitchen scene with focus on the microwave and oven.
6
An empty wooden bench near a large body of water.
7
A brown bear on two legs is near the water.
8
An adult and a baby elephant crossing a road.
9
A man rides a motorcycle on a track.
10
A bear that is in the grass in the wild.
11
A black-and-white photo of a large bench on the sidewalk.
12
A view of a sink and toilet in a bathroom in the dark.

### v3

@@ 2.2 Differential Privacy
Let $\text{SG}_{q,\sigma}(D)$ be defined as Definition 2.5 for function $f$. Then $\text{SG}_{q,\sigma}(D)$ satisfies $(\alpha,\epsilon)$-RDP with
@@ 4.2 Quantitative Results
We first demonstrate the performance of DP-RDM quantitatively in terms of the Fréchet inception distance (FID) score. Figure 6 shows the FID score for DP-RDM on MS-COCO and Shutterstock when answering $T$ queries privately for $T$ ranging from 1 to 10k. The evaluation dataset and the private retrieval dataset are the same, i.e. for Figure 6(a) we treat the MS-COCO validation set as the evaluation dataset and MS-COCO training set as the private retrieval dataset. We make the following observations.
@@ 4.3 Qualitative Results
Ref Validation Prompt
1 image of blue sky with white cloud for background usage.
2 wood texture paint yellow
3 School of Jackfish in Sipadan Malaysia
4 A bunch of garlic isolated on white background
5 Halloween pumpkins on wooden table on dark color background
6 cauliflower
7 Closeup of golden art sculpture pagoda and golden tiered umbrella in temple against with blue sky in north of Thailand. Faith and religion concept.
8 White stork hunting in a meadow
9 abstract light background
10 White grunge wood close-up background. Painted old wooden wall.
11 Multicolored abstract background holidays lights in motion blur image. illustration digital.
12 Torn blue jean fabric texture.
@@ 8 White stork hunting in a meadow
Ref Validation Prompt
1 A dog watching a television that is displaying a picture of a dog.
2 The Asian girls are sharing conversation under a parasol.
3 An aerial view of several jumbled cars on a narrow road.
4 a short passenger train on a track by a grassy hill
5 A kitchen scene with focus on the microwave and oven.
6 An empty wooden bench near a large body of water.
7 A brown bear on two legs is near the water.
8 An adult and a baby elephant crossing a road.
9 A man rides a motorcycle on a track.
10 A bear that is in the grass in the wild.
11 A black-and-white photo of a large bench on the sidewalk.
12 A view of a sink and toilet in a bathroom in the dark.